# The Behrens–Fisher problem revisited

### v1

@@ 3.1 The equal-variance $(g=h)$ case
which is the standard Studentt-t density with $\nu=n_{1}+n_{2}-2$ degrees of freedom. The equal-variance case clarifies how the $t$-scale $1+t^{2}/\nu$ emerges inside the more general factor $\left[1+t^{2}/2\right]$. This is crucial when we later study saddle-point expansions: any correct asymptotic must hit this benchmark as $g\to h$. When a library routine for $_{2}F_{1}$ is called, the very first check in robust code is: “if $|\rho|<\varepsilon$ $\Rightarrow$ use Studentt-t”. This avoids catastrophic loss of accuracy from subtractive cancellation when $g$ and $h$ differ only in the 10th decimal place, thereby guiding numerical implementation.
@@ 4 Ramanujan’s master theorem for tail analysis
For any even density $f(t)$, the algebraic tail $f(t)\sim\sum_{m=0}^{\infty}(-1)^{m}A_{m}|t|^{-(2m+1)}$ is precisely the inverse Mellin transform of $\mathcal{M}{f}(s)=\int_{0}^{\infty}t^{s-1}f(t)dt$, taken at negative odd integers s=2m1s=-2m-1. Hence, if we possess a MB representation of $f$, Ramanujan’s master theorem lets us jump straight to the numeric coefficients $A_{m}$ without evaluating any integrals. That is invaluable for building Edgeworth or Cornish–Fisher corrections; deriving saddle-point approximations; and bounding Type-I error when $\lvert t\rvert$ exceeds conventional tables. In our work, we had the one-fold contour $\mathcal{M}{f_{T}}(s)=\frac{1}{2\pi i}\int\limits_{c-i\infty}^{c+i\infty}\Phi(z)K(z,s)dz$, where $\Phi(z)=\Gamma(-z)\Gamma\left(\tfrac{1}{2}+z\right)\Gamma\left(z+\tfrac{\nu_{1}}{2}\right)\Gamma\left(\tfrac{\nu_{2}}{2}-z+\tfrac{1}{2}\right)$ and $K(z,s)$ is a benign product of powers $g,h,\alpha]$ that is analytic in $z$ near $z=0$. All the pole structure is locked inside the $\Gamma$-product $\Phi(z)$. Set $A(s)=\Phi\left(-\tfrac{s+1}{2}\right)$, $s\in(-1-\nu_{2},\nu_{1})$, so that the Mellin transform around s=1s=-1 (odd negative integers) is governed by $A(s)$. Divide A(s)A(-s) by $\Gamma(s)$ and expand at $s=0$ using the rising factorial series of $\Gamma$ near the origin to get $\frac{A(-s)}{\Gamma(s)}=\sum_{m=0}^{\infty}\frac{(-1)^{m}}{m!}a_{m}s^{m}$, where $a_{m}=\frac{\Gamma\left(m+\tfrac{1}{2}\right)\Gamma\left(m+\tfrac{\nu_{1}}{2}\right)\Gamma\left(m+\tfrac{\nu_{2}}{2}\right)}{\sqrt{\pi}}$. Ramanujan’s master theorem identifies, for $m=0,1,\dots$,
$A_{m}$ $=$ $\frac{\Gamma\left(m+\tfrac{1}{2}\right)\Gamma\left(m+\tfrac{\nu_{1}}{2}\right)\Gamma\left(m+\tfrac{\nu_{2}}{2}\right)}{\sqrt{\pi}2^{\nu_{1}/2+\nu_{2}/2+m}(g/\nu_{1})^{\nu_{1}/2}(h/\nu_{2})^{\nu_{2}/2}},$ (31)
@@
where $A_{m}$ is given by (31).
The leading term gives
@@
Ramanujan’s theorem leads to (i) Uniform large-deviation approximations, by combining (32) with an exponential tilt (Cramér transform described by [30, Chapter 2]) to construct uniform–in–$t$ approximations akin to Daniels’ formula (see [31]) for the signed–root $\chi^{2}$; (ii) Moment recovery: the Mellin transform at positive integers $s=1,3,5,\ldots$ gives absolute moments $\mathbb{E}\lvert T\rvert^{s-1}$. Ramanujan’s coefficients analytically continue those moments via the functional equation of $\Gamma$, bypassing direct integration; (iii) Edgeworth correction for the cdf, by integrating (32) term-wise yields a descending series in $\lvert t\rvert^{-2m}$ for $1-F_{T}(t)$. Two terms often slash the upper–tail $p$-value error below $10^{-4}$; (iv) Automatic symbolic framework: since $A_{m}$ is an algebraic function of $\nu_{1},\nu_{2}$ and $g/h$, computer algebra can emit tail expansions on-demand for arbitrary parameter triples. Ramanujan’s master theorem serves as an analytical lens, revealing—through the MB integral—the precise asymptotic coefficients that dictate the decay rate and smoothness of the Behrens–Fisher density at infinity. These coefficients, which are otherwise obscured within layers of hypergeometric complexity, become immediately accessible via this transform-based approach. Once obtained, they enable the analyst to craft sharp $p$-value approximations, test practical significance where Monte Carlo is slow, and systematically evaluate any heuristic variance-unequal-$t$-procedure by benchmarking it against the exact analytic benchmark provided by the hypergeometric representation.
@@ 5 Discussion
$\nu_{1}$ 113030 in steps of $1$, then $40$, $60$, $120$ Covers classical small–sample and asymptotics
$\nu_{2}$ Same as $\nu_{1}$ Rectangular grid; use symmetry to halve storage
$r$ $\{0.1,0.2,\dotsc,1.0\}$; refine on log scale near 1 (e.g., 0.85, 0.9, 0.95) Finer mesh where Welch performs worst
@@
$8<t\leq 15$ Ramanujan’s master theorem tail series Summed until term $10^{-16}$. Only 3-5 terms suffice; avoids cancellation near zero-crossings.
$t>15$ Saddle-point approximation Uses Ramanujan’s master theorem-based remainder bounds. Achieves relative error $10^{-14}$.
@@ 5.2 Edge-case tail error using Lugannani–Rice saddle-points
We now develop the Lugannani–Rice (LR) saddle-point formula and its implementation for the Behrens–Fisher statistic. Using this, we identify the edge-cases, i.e., parameter corners where LR may exhibit anomalous or unstable behavior. We will explore how to use the exact Ramanujan series coefficients $A_{m}$ as a reference yardstick for tail error, alongside some practical rules and diagnostic plots. From the MB contour we have, for $\lvert\xi\rvert<\nu_{2}$,
@@
Region Potential issue Symptom Detail
Tiny d.f. $\nu_{1}\lor\nu_{2}\leq 2$ DiGamma behaves like $\tfrac{2}{\nu}+\dots$; Newton steps may leave domain Non-convergence or absurd $\hat{\xi}$ Extreme curvature causes oscillation or exit from feasible range
Extreme variance ratio $r\ll 0.1$ or $r\gg 10$ Saddle lies near domain edge: $\xi\to\nu_{2}^{-}$ or $\xi\to-\nu_{1}^{+}$ $w\to 0$; inflated LR error Numerical instability from vanishing weights near boundary
Ultra-high tail $t>0.8(\nu_{1}+\nu_{2})$ Third cumulant diverges; LR loses $o(t^{-3})$ accuracy Error exceeds $10^{-3}$ even for moderate d.f. Saddle-point approximation fails due to excess skewness
Near equal variance but small $n$ ($r\approx 1$, $\nu_{i}\leq 4$) Cancellation in $K_{T}^{\prime}(\xi)$ causes floating-point loss LR discontinuity vs exact curve Significant digit loss appears as sharp artifact in tail plots
Table 5: Parameter regimes where the LR saddle-point method may exhibit instability.
@@
To rigorously test the implementation and robustness of the LR tail approximation, we construct a comprehensive evaluation grid over a representative range of parameters. Specifically, degrees of freedom are selected as $\nu_{i}\in\{1,2,5,10,20,40\}$ for each sample, while the variance ratio grid spans $r=g/h\in\{0.05,0.1,0.25,0.5,1,2,4,10,20\}$. Tail probabilities are evaluated at test statistics $t\in\{6,8,10,12,15,20\}$, probing both moderate and extreme regions of the distribution. The exact reference tail probability is computed using two methods depending on the magnitude of $t$. For moderate values ($t\leq 8$), the integral representation involving the Appell function is numerically evaluated using adaptive quadrature (e.g., QUADPACK). For larger $t$, where the integral becomes numerically unstable, Ramanujan’s master theorem yields a rapidly converging tail series. The LR approximation can be implemented with all calculations performed in extended precision arithmetic to avoid catastrophic cancellation in the diGamma difference terms. As a safeguard, we flag a trial as non-convergent if the Newton iteration produces a solution violating $\lvert\hat{\xi}\rvert>\nu_{\min}-10^{-7}$, indicating a step outside the feasible domain. To quantify performance, the relative error in tail probability is computed via the metric
@@
$\delta_{\text{rel}}(t)=\frac{\lvert\Pr_{\text{LR}}(T\geq t)-\Pr_{\text{exact}}(T\geq t)\rvert}{\Pr_{\text{exact}}(T\geq t)}.$ (35)
This error is visualized using log–scaled heatmap, with $\log_{10}\delta_{\text{rel}}(t)$ plotted against $t$ on the horizontal axis and $\log_{10}r$ on the vertical axis, separately for each $(\nu_{1},\nu_{2})$ configuration. These visual summaries help identify parameter regimes where the LR approximation remains accurate, and where it may break down.
Several qualitative patterns emerge from the error landscape. First, there exists a broad band of numerical reliability: for moderate variance ratios satisfying $|\log_{10}r|\leq 1$ and degrees of freedom $\nu_{i}\geq 5$, the LR approximation maintains high accuracy, with relative error $\delta_{\text{rel}}<10^{-4}$ uniformly up to $t=12$. In contrast, variance-dominated tails present a challenge: when $r<0.1$ and $t>0.7\nu_{2}$, the LR error exceeds 1%, as the saddle point approaches the boundary of the parameter domain where the density is exponentially suppressed. In such regimes, adding just one more term from the Ramanujan expansion (e.g., $m=2$) halves the discrepancy, but the LR approximation still remains an order of magnitude off. Finally, in the low degrees of freedom regime, particularly when $\nu_{1}=1$ or $2$ and $t>8$, Newton’s method tends to overshoot the saddle unless damped—e.g., by applying a step factor of 0.5. After damping, the approximation yields $\delta_{\text{rel}}\approx 5\times 10^{-3}$, which, while not high-precision, remains adequate for rough $p$-value estimation.
@@
In small-to-moderate samples the Welch test calibrates its critical value through the random denominator $\nu_{W}$ in an attempt to mimic the heavier tails of the true Behrens–Fisher distribution, whose algebraic decay rate is fixed by the total degrees of freedom $\nu_{1}+\nu_{2}$. When the sample–variance ratio $r=\sigma_{1}^{2}/\sigma_{2}^{2}$ is near unity, $\nu_{W}$ fluctuates tightly around $\nu_{1}+\nu_{2}$, so Welch’s reference distribution matches the exact tail probability and the test is essentially exact. As $r$ drifts away from 1, however, the larger sample variance increasingly dominates the pooled standard error while simultaneously shrinking its own contribution to the denominator of $\nu_{W}$; this forces $\nu_{W}$ below $\nu_{1}+\nu_{2}$, producing a $t$-law with thicker tails than the truth and hence a liberal Type-I error. Conversely, if the smaller variance attaches to the smaller sample, $\nu_{W}$ overshoots $\nu_{1}+\nu_{2}$, the reference tails become too light, and the test turns conservative. The sign-flip contour seen in the heat-maps is precisely the parameter locus where the expectation of $\nu_{W}$ crosses $\nu_{1}+\nu_{2}$; on that curve the tail exponents balance, and Welch’s nominal size coincides with the exact level.
5.3 Welch versus compact Behrens–Fisher
@@
To assess the finite-sample performance of the Welch approximation relative to the exact Behrens–Fisher solution, we examine the deviation $\Delta(n,r)=\Pr_{\text{exact}}\left(\lvert T\rvert>t_{\nu_{W},0.025}\right)-0.05$, where $n_{1}=n_{2}=n$ and $r=\sigma_{1}^{2}/\sigma_{2}^{2}$ is the population variance ratio. This quantity measures the size distortion introduced by Welch’s test when using the standard 5% critical value from the $t_{\nu_{W}}$ distribution. The values of $\Delta(n,r)$ were estimated using Monte Carlo simulation with 5000 replications per grid point and plotted as a heatmap shown in Fig. 1. Rows correspond to sample sizes $n\in\{10,20,30,50\}$, and columns to $\log_{2}r\in\{-3,-2,-1,0,1,2,3\}$. The black contour in the figure marks the zero crossing $\Delta(n,r)=0$, i.e., the crossover locus at which Welch’s test shifts from being conservative (size below 5%) to liberal (size above 5%). Blue regions indicate conservative behavior; red regions indicate liberal behavior.
A summary of where these sign flips occur is provided in Table 6. This behavior reveals that for smaller sample sizes (e.g., $n\leq 20$), Welch’s test becomes liberal as soon as one sample variance exceeds the other by a factor of about 2, and conservative when the reverse ratio exceeds roughly 2.5 to 3. In contrast, for larger samples (e.g., $n=50$), the safe range of variance ratios—where the size remains close to nominal—widens considerably. At $n=50$, Welch remains within $\pm 0.2\%$ of the nominal size as long as $0.6\lesssim r\lesssim 2.8$. These results yield a practical rule of thumb: when $n_{1}=n_{2}\leq 30$, flag potential size distortion whenever the sample variance ratio exceeds 2:1 in either direction. Such imbalances can materially affect the Type I error rate under Welch’s approximation, especially in hypothesis testing contexts with small to moderate sample sizes. The crossover contour $\Delta(n,r)=0$ can be tabulated to support automated test selection protocols. For example, statistical software can dynamically switch from Welch’s $t$ to an exact or LR–based critical value whenever the observed $(n,r)$ lies outside a pre-defined “safe region.” For future work, replacing the Monte Carlo–based exact tail probability with the analytic approximation—e.g., via Lugannani–Rice coupled with a Ramanujan’s master theorem tail expansion—offers a pathway toward high–resolution inference for the $n_{1}\neq n_{2}$ case. An analytical comparison between Welch’s approximation and the work presented in this paper is summarized in Table 7.
@@
Property Welch’s test Present work
Representation Student-$t_{\nu_{W}}$ with random degrees of freedom Deterministic hypergeometric pdf/cdf
Exactness First-order accurate as $n\to\infty$ Exact for all $(n_{i},r)$
Symmetry under $(\nu_{1},\nu_{2},r)\leftrightarrow(\nu_{2},\nu_{1},1/r)$ Broken Preserved
Tail order $t^{-(\nu_{W}+1)}$ $t^{-(\nu_{1}+\nu_{2}+1)}$ (correct)
Error control No analytic bound; relies on simulation Ramanujan’s master theorem coefficients yield explicit truncation error: $O(t^{-5})$
Computational cost (single $p$-value) $O(1)$ $O(1)$ (one $_{2}F_{1}$ call or $\leq 10$ residue terms)
Table 7: Comparison of Welch’s approximation and the present work.
@@
5.4 Nel et al. [28] versus compact Behrens–Fisher
Aspect Nel et al [28] Our work
Univariate Behrens–Fisher pdf Yes: one Gauss’ hypergeometric expression Yes: two algebraically equivalent forms—(i) a terminating/convergent residue series; (ii) a hypergeometric representation
Univariate cdf Not provided Provided: a one-dimensional Euler–Beta integral reducible to the Appell $F_{1}$; closed formula valid for all $(n_{1},n_{2})$
Tail expansion & error bounds No Ramanujan-based $t^{-(2m+1)}$ coefficients with computable remainder
Saddle-point diagnostics No Lugannani–Rice with explicit cumulants from the MB generator; uniform $O(t^{-4})$ error away from edges
Critical-value tables No Exact 0.10–0.001 quantiles for any $(n_{i},r)$ ( a sample CSV file in supplement)
Welch benchmarking No Heatmap and crossover contour locating the sign flip of Welch’s test
Multivariate extension Yes (outline for Hotelling $T^{2}$) Deferred to future work; univariate derivations generalizable
Table 8: Comparison between the contributions of Nel et al. [28] and our work.

### v2

@@ 3.1 The equal-variance $(g=h)$ case
the standard Studentt-t density with $\nu=n_{1}+n_{2}-2$ degrees of freedom. The $g=h$ case clarifies how the $t$-scale $1+t^{2}/\nu$ emerges inside the more general factor $\left[1+t^{2}/2\right]$. This is crucial when we study saddle-point expansions: any correct asymptotic must hit this benchmark as $g\to h$. In library routines, the first check is “if $|\rho|<\varepsilon$ $\Rightarrow$ use Studentt-t”, avoiding loss of accuracy when $g$ and $h$ differ marginally, thereby guiding numerical implementation.
@@ 4 Ramanujan’s master theorem for tail analysis
For any even density $f(t)$, the algebraic tail $f(t)\sim\sum_{m=0}^{\infty}(-1)^{m}A_{m}|t|^{-(2m+1)}$ is precisely the inverse Mellin transform of $\mathcal{M}{f}(s)=\int_{0}^{\infty}t^{s-1}f(t)dt$, taken at negative odd integers s=2m1s=-2m-1. Hence, if we possess a MB representation of $f$, RMT leads straight to the numeric coefficients $A_{m}$ without evaluating any integrals—useful for Edgeworth corrections; saddle-point approximations; and bounding Type-I error when $\lvert t\rvert$ exceeds conventional tables. In our work, we had the one-fold contour $\mathcal{M}{f_{T}}(s)=\frac{1}{2\pi i}\int\limits_{c-i\infty}^{c+i\infty}\Phi(z)K(z,s)dz$, where $\Phi(z)=\Gamma(-z)\Gamma\left(\tfrac{1}{2}+z\right)\Gamma\left(z+\tfrac{\nu_{1}}{2}\right)\Gamma\left(\tfrac{\nu_{2}}{2}-z+\tfrac{1}{2}\right)$ and $K(z,s)$ is a benign product of powers $g,h,\alpha]$ that is analytic in $z$ near $z=0$. All the pole structure is locked inside the $\Gamma$-product $\Phi(z)$. Set $A(s)=\Phi\left(-\tfrac{s+1}{2}\right)$, $s\in(-1-\nu_{2},\nu_{1})$, so that the Mellin transform around s=1s=-1 (odd negative integers) is governed by $A(s)$. Divide A(s)A(-s) by $\Gamma(s)$ and expand at $s=0$ using the rising factorial series of $\Gamma$ near the origin to get $\frac{A(-s)}{\Gamma(s)}=\sum_{m=0}^{\infty}\frac{(-1)^{m}}{m!}a_{m}s^{m}$, where $a_{m}=\frac{\Gamma\left(m+\tfrac{1}{2}\right)\Gamma\left(m+\tfrac{\nu_{1}}{2}\right)\Gamma\left(m+\tfrac{\nu_{2}}{2}\right)}{\sqrt{\pi}}$. RMT identifies
$A_{m}$ $=$ $\frac{\Gamma\left(m+\tfrac{1}{2}\right)\Gamma\left(m+\tfrac{\nu_{1}}{2}\right)\Gamma\left(m+\tfrac{\nu_{2}}{2}\right)}{\sqrt{\pi}2^{\nu_{1}/2+\nu_{2}/2+m}(g/\nu_{1})^{\nu_{1}/2}(h/\nu_{2})^{\nu_{2}/2}},\quad m=0,1,\dots,$ (18)
@@
where $A_{m}$ is given by (18). The leading term gives
@@
RMT leads to (i) uniform large-deviation approximations, by combining (19) with an exponential tilt (Cramér transform described by (Dembo and Zeitouni, 1998, Chapter 2)) to construct uniform–in–$t$ approximations akin to Daniels’ formula (see Daniels (1954)) for the signed–root $\chi^{2}$; (ii) moment recovery: the Mellin transform at positive integers $s=1,3,5,\ldots$ gives absolute moments $\mathbb{E}\lvert T\rvert^{s-1}$. Ramanujan’s coefficients analytically continue those moments via the functional equation of $\Gamma$, bypassing direct integration; (iii) Edgeworth correction for the cdf, by integrating (19) term-wise yields a descending series in $\lvert t\rvert^{-2m}$ for $1-F_{T}(t)$. Two terms often slash the upper–tail $p$-value error below $10^{-4}$; (iv) automatic symbolic framework: since $A_{m}$ is an algebraic function of $\nu_{1},\nu_{2}$ and $g/h$, computer algebra can emit tail expansions on-demand for arbitrary parameter triples. RMT reveals—through the MB integral—the precise asymptotic coefficients that dictate the decay rate and smoothness of the Behrens–Fisher density at infinity. These coefficients, which are otherwise obscured within layers of hypergeometric complexity, become immediately accessible via this transform-based approach. Once obtained, they enable the analyst to craft sharp $p$-value approximations, test practical significance where Monte Carlo is slow, and systematically evaluate any heuristic variance-unequal-$t$-procedure by benchmarking it against the exact analytic benchmark provided by the hypergeometric representation.
In the next section, we provide detailed discussions including specifications for cdf tables, LR saddlepoint approximations, and comparative analyses of our results with those of Welch (1947) and Nel et al. (1990).
@@ 5 Discussion
$\nu_{1}$ 113030 in steps of $1$, then $40$, $60$, $120$ Covers classical small–sample and asymptotics.
$\nu_{2}$ Same as $\nu_{1}$ Rectangular grid; use symmetry to halve storage.
$r$ $\{0.1,0.2,\dotsc,1.0\}$; refine on log scale near 1 Finer mesh where Welch performs worst.
@@
$8<t\leq 15$ RMT tail series Summed until term $10^{-16}$. Only 3-5 terms suffice; avoids cancellation near zero-crossings.
$t>15$ Saddle-point approximation Uses RMT-based remainder bounds. Achieves relative error $10^{-14}$.
@@ 5.2 Edge-case tail error using Lugannani–Rice saddle-points
We now develop the LR saddle-point formula and its implementation for the Behrens–Fisher statistic. Using this, we identify the edge-cases, i.e., parameter corners where LR may exhibit anomalous or unstable behavior. We will explore how to use the exact Ramanujan series coefficients $A_{m}$ as a reference yardstick for tail error, alongside some practical rules and diagnostic plots. From the MB contour we have, for $\lvert\xi\rvert<\nu_{2}$,
@@
Figure 1: LR tail approximation against the exact hypergeometric reference.
We evaluated the LR tail approximation against the exact hypergeometric reference, for representative degrees of freedom pairs $(\nu_{1},\nu_{2})\in\{(5,5),(10,10),(2,10),(20,5)\}$, by computing $P(T\geq t)$ over a grid spanning $r\in\{0.05,0.1,0.25,0.5,1,2,4,10,20\}$ and $t\in\{6,8,10,12,15,20\}$. The exact tail was obtained by numerically integrating our closed-form hypergeometric density from $t$ to $\infty$ using an adaptive split with an exponential change of variables; the LR approximation used the cumulant generator and Newton’s method for the saddlepoint in extended precision. We visualized accuracy via
@@
$\log_{10}\delta_{\mathrm{rel}}(t)=\log_{10}\left(\frac{\lvert\Pr_{\text{LR}}(T\geq t)-\Pr_{\text{exact}}(T\geq t)\rvert}{\Pr_{\text{exact}}(T\geq t)}\right)$
heat-maps. Several qualitative patterns emerge (see Fig. 1) from the numerical error analysis. Across balanced or moderately imbalanced variances ($|\log_{10}r|\leq 1$) and moderate tails, LR errors were uniformly small (typically $\leq 10^{-3}$). As predicted, accuracy degrades in variance-extreme corners and very far tails, where the saddle approaches the domain boundary; these regions exhibit elevated $\delta_{\mathrm{rel}}$. In such regimes, adding just one more term from the Ramanujan expansion (e.g., $m=2$) halves the discrepancy, but the LR approximation still remains an order of magnitude off. In the low degrees of freedom regime, particularly when $\nu_{1}=1$ or $2$ and $t>8$, these effects are amplified and sometimes require damping in the Newton iteration. Overall, the plots delineate a broad reliability band for LR and clearly mark the edge regimes where the Ramanujan expansion should be preferred. These visual summaries help identify parameter regimes where the LR approximation remains accurate, and where it may break down.
The LR saddle-point method is reliable over most of the grid but has recognizable edge cases. When either degrees of freedom is tiny $\nu_{1}\leq 2$ or $\nu_{2}\leq 2$, the digamma term behaves like $2/\nu+\cdots$, so Newton updates for the saddle $\hat{\xi}$ can step outside the admissible domain; the optimizer then oscillates or returns implausible $\hat{\xi}$. For extreme variance ratios ($r\ll 0.1$ or $r\gg 10$), the saddle approaches the boundary $\xi\to\nu_{2}^{-}$ or $\xi\to-\nu_{1}^{+}$, driving $w\to 0$; LR’s error inflates and numerical weights nearly vanish, producing instability. In ultra-high tails $t>0.8(\nu_{1}+\nu_{2})$, the third cumulant grows and LR loses its $o(t^{-3})$ accuracy; absolute errors can exceed $10^{-3}$ even at moderate d.f. Finally, when variances are nearly equal but sample sizes are small ($r\approx 1$, $\nu_{i}\leq 4$), cancellations in $K_{T}^{\prime}(\xi)$ cause significant-digit loss, seen as small discontinuities between LR and the exact curve in tail plots. Using results from RMT, we get $\Pr(T\geq t)=\sum_{m=0}^{M-1}\frac{(-1)^{m+1}A_{m}}{m!\,(2m)}\,t^{-2m}+R_{M}(t)$, where $A_{m}$ is given by (18). For $\lvert t\rvert\geq 10$ the first two terms ($m=0,1$) give $10^{-10}$ relative error, so we treat $F_{\text{RMT}}(t)=\frac{A_{0}}{t}-\frac{A_{1}}{2t^{3}}$ as ground truth in the far tail.
@@
In small-to-moderate samples, the Welch test calibrates its critical value through the random denominator $\nu_{W}$ in an attempt to mimic the heavier tails of the true Behrens–Fisher distribution, whose algebraic decay rate is fixed by the total degrees of freedom $\nu_{1}+\nu_{2}$. When the ratio $r=\sigma_{1}^{2}/\sigma_{2}^{2}$ is near unity, $\nu_{W}$ fluctuates tightly around $\nu_{1}+\nu_{2}$, so Welch’s reference distribution matches the exact tail probability and the test is essentially exact. As $r$ drifts away from 1, however, the larger sample variance increasingly dominates the pooled standard error while simultaneously shrinking its own contribution to the denominator of $\nu_{W}$; this forces $\nu_{W}$ below $\nu_{1}+\nu_{2}$, producing a $t$-law with thicker tails than the truth and hence a liberal Type-I error. Conversely, if the smaller variance attaches to the smaller sample, $\nu_{W}$ overshoots $\nu_{1}+\nu_{2}$, the reference tails become too light, and the test turns conservative. The sign-flip contour seen in the heat-maps in Section 5.3 is precisely the parameter locus where the expectation of $\nu_{W}$ crosses $\nu_{1}+\nu_{2}$; on that curve the tail exponents balance, and Welch’s nominal size coincides with the exact level.
5.3 Welch’s approximation versus our unified approach
@@
To assess the finite-sample performance of the Welch approximation relative to the exact Behrens–Fisher solution, we examine the deviation $\Delta(n,r)=\Pr_{\text{exact}}\left(\lvert T\rvert>t_{\nu_{W},0.025}\right)-0.05$, where $n_{1}=n_{2}=n$ and $r=\sigma_{1}^{2}/\sigma_{2}^{2}$ is the population variance ratio. This quantity measures the size distortion introduced by Welch’s test when using the standard 5% critical value from the $t_{\nu_{W}}$ distribution. The values of $\Delta(n,r)$ were estimated using Monte Carlo simulation with 5000 replications per grid point and plotted as a heatmap shown in Fig. 2. Rows correspond to sample sizes $n$, and columns to $\log_{2}r\in\{-3,-2,-1,0,1,2,3\}$. The black contour in the figure marks the zero crossing $\Delta(n,r)=0$, i.e., the crossover locus at which Welch’s test shifts from being conservative (size below 5%) to liberal (size above 5%). Blue regions indicate conservative behavior; red regions indicate liberal behavior.
A summary of where these sign flips occur is provided in Table 5. This behavior reveals that for smaller sample sizes (e.g., $n\leq 20$), Welch’s test becomes liberal as soon as one sample variance exceeds the other by a factor of about 2, and conservative when the reverse ratio exceeds roughly 2.5 to 3. In contrast, for larger samples (e.g., $n=50$), the safe range of variance ratios—where the size remains close to nominal—widens considerably. At $n=50$, Welch remains within $\pm 0.2\%$ of the nominal size as long as $0.6\lesssim r\lesssim 2.8$. These results yield a practical rule of thumb: when $n_{1}=n_{2}\leq 30$, flag potential size distortion whenever the sample variance ratio exceeds 2:1 in either direction. Such imbalances can materially affect the Type I error rate under Welch’s approximation, especially in hypothesis testing contexts with small to moderate sample sizes. The crossover contour $\Delta(n,r)=0$ can be tabulated to support automated test selection protocols. For example, statistical software can dynamically switch from Welch’s $t$ to an exact or LR–based critical value whenever the observed $(n,r)$ lies outside a pre-defined “safe region.” For future work, replacing the Monte Carlo–based exact tail probability with the analytic approximation—e.g., via Lugannani–Rice coupled with a Ramanujan’s master theorem tail expansion—offers a pathway toward high–resolution inference for the $n_{1}\neq n_{2}$ case. An analytical comparison between Welch’s approximation and our work is presented in the following paragraph.
@@
Welch’s test models the statistic with a Student distribution having a random degrees of freedom $\nu_{W}$—yielding a first-order-accurate approximation as $n\to\infty$—but it breaks the natural symmetry $(\nu_{1},\nu_{2},r)\leftrightarrow(\nu_{2},\nu_{1},1/r)$. Its tail decays like $t^{-(\nu_{W}+1)}$, and there is no analytic error bound—assessment typically relies on simulation. By contrast, our approach provides a deterministic closed form for pdf/cdf via a single Gauss hypergeometric function (or, a residue series), is exact for every $(n_{1},n_{2},r)$, and preserves the symmetry under index/ratio inversion. The tail order is the correct $t^{-(\nu_{1}+\nu_{2}+1)}$. Moreover, RMT gives explicit inverse-power coefficients, so truncation errors are certified (e.g., $O(t^{-5})$ after two correction terms), furnishing analytic control absent in Welch. Computationally, both yield $O(1)$ $p$-values: Welch evaluates a $t$-cdf, while our method requires one $_{2}F_{1}$ call (or, at most $\leq 10$ residue terms) and is comparably fast in practice.
@@
5.4 Nel et al. (1990) versus our work